OpenStudy (anonymous):

solve an airplain flew for 4 hours against a head wind of 40km/h. on the return flight the same wind was now a tail wind and the fight took 3 hours. find the speed of the airplain in sill air

OpenStudy (anonymous):

First lets assume that the velocity of the airplane is v

OpenStudy (anonymous):

And also we assume that the velocity of the wind is w

OpenStudy (anonymous):

S= d/t s - 40 = d /4 S + 40 = d / 3

OpenStudy (anonymous):

(v-w)4=(v+w)3

OpenStudy (anonymous):

D = 4s -160 D = 3s + 120 therefore, 4s -160 = 3s + 120 s = 120 + 160 s = 280 km / hr

OpenStudy (anonymous):

Now in place of w just put the value 40 and solve the equation for v